What is the largest possible area of a rectangle with a perimeter of 500 feet?
Answer by checkley77(12844)

500/4=125 FT. IS THE SIDE MEASUREMENT OF THE SQUARE WHICH IS THE MAXIMUM AREA.
125^2=15,625 FT^2. MAX AREA.
PROOF:
124*126=15,624 FT^2.